OpenStudy (anonymous):

A recipe calls for 3 cans of tomato sauce. Fawn wants to make 40% fewer servings than the recipe makes. Which expression represents how many cans of tomato sauce Fawn should use?

OpenStudy (anonymous):

http://prntscr.com/72xu1m

OpenStudy (anonymous):

@KitKat16

OpenStudy (anonymous):

If Fawn wants to make 40% less, that means they want to make 60% of the original. So just choose the one that would be \[.6*3\]

OpenStudy (anonymous):

18

OpenStudy (anonymous):

No that's a .6 not a 6. And don't solve for it, choose the one that is equal to it out of the choices.

OpenStudy (anonymous):

I dont get what you saying

OpenStudy (anonymous):

What fraction is equal to .6?

OpenStudy (anonymous):

2 3

OpenStudy (kitkat16):

60%*3 = 9/5

OpenStudy (anonymous):

what? What would the decimal, .6, be in fraction form?

OpenStudy (anonymous):

0.6

OpenStudy (kitkat16):

no find the equation for 9/5

OpenStudy (kitkat16):

math isn't my best subject I don't want to tell you the wrong steps

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

.6 is a decimal, what would it be as a fraction? As in a number over a number.

OpenStudy (kitkat16):

@LegendarySadist I thought 9/5 =3/5*3

OpenStudy (anonymous):

0.6 i said

OpenStudy (anonymous):

yes, that would be it, but .6 = 3/5 and that's what I'm trying to get OP to understand.

OpenStudy (kitkat16):

ok sorry

OpenStudy (anonymous):

0.6 is the same as .6. Those are both decimals. A fraction is a number over a number. What is 0.6 in fraction form?

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

Yes, it is B.

OpenStudy (anonymous):

I PASSED!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Congrats :)

OpenStudy (kitkat16):

:)